OpenStudy (ellie202000):

Y+3=5/2(x+13) I have no idea what I am doing wrong. I keep getting y=5/2x+11 but that's not right.

OpenStudy (solomonzelman):

\(\displaystyle y+3=\frac{5}{2}(x+13)\) like this?

OpenStudy (ellie202000):

yes

OpenStudy (solomonzelman):

Above is the *point slope* form of the equation, and you (as I see) would like to convert this to a *slope intercept* form.

OpenStudy (solomonzelman):

\(\displaystyle y+3=\frac{5}{2}(x+13)\) \((1)\) First, you have to expand the right side. (Knowing that: *a*(b+c)= *a*b + *a*c) \((2)\) Then, subtract 3 from both sides. (And there will be your answer)

OpenStudy (ellie202000):

So what step am I missing? \[y+3=\frac{ 5 }{ 2 }(x+13)\] \[y+3-3=\frac{ 5 }{ 2 }x+14-3\] \[y=\frac{ 5 }{ 2 }x-11\]

OpenStudy (ellie202000):

@pooja195 @mathstudent55 @ParthKohli @SolomonZelman

OpenStudy (ellie202000):

The answer is \[y=\frac{ 5}{ 2 }x+\frac{ 59 }{ 2 }\]